# Differentiable Task Graph Learning: Procedural Activity Representation and Online Mistake Detection from Egocentric Videos

### v1

@@ Abstract
Procedural activities are sequences of key-steps aimed at achieving specific goals. They are crucial to build intelligent agents able to assist users effectively. In this context, task graphs have emerged as a human-understandable representation of procedural activities, encoding a partial ordering over the key-steps. While previous works generally relied on hand-crafted procedures to extract task graphs from videos, in this paper, we propose an approach based on direct maximum likelihood optimization of edges’ weights, which allows gradient-based learning of task graphs and can be naturally plugged into neural network architectures. Experiments on the CaptainCook4D dataset demonstrate the ability of our approach to predict accurate task graphs from the observation of action sequences, with an improvement of +16.7% over previous approaches. Owing to the differentiability of the proposed framework, we also introduce a feature-based approach, aiming to predict task graphs from key-step textual or video embeddings, for which we observe emerging video understanding abilities. Task graphs learned with our approach are also shown to significantly enhance online mistake detection in procedural egocentric videos, achieving notable gains of +19.8% and +7.5% on the Assembly101 and EPIC-Tent datasets. Code for replicating experiments is available at https://github.com/fpv-iplab/Differentiable-Task-Graph-Learning.
@@ 1 Introduction
Previous approaches focused on directly tackling tasks requiring procedural knowledge such as action anticipation [15, 13, 30] and mistake detection [33, 12, 6, 37, 14] without developing explicit representations of the procedure. Other works proposed neural models able to develop implicit representations of the procedure by learning how to recover missing actions [39, 25], discover key-steps [10, 4, 5], or grounding them to video [9, 22]. A different approach [3, 9, 16] consists in representing the structure of a procedure in the form of a task graph, i.e., a Directed Acyclic Graph (DAG) in which nodes represent key-steps, and directed edges impose a partial ordering over key-steps, encoding dependencies between them (see Figure 1(a)). Graphs provide an explicit representation which is readily interpretable by humans and easy to incorporate in downstream tasks such as detecting mistakes or validating the execution of a procedure. While graphs have been historically used to represent constraints in complex tasks and design optimal sub-tasks scheduling [34], graph-based representations mined from videos [3], key-step sequences [35, 18] or external knowledge bases [40] have only recently emerged as a powerful representation of procedural activities able to support downstream tasks such as key-step recognition or forecasting [3, 40]. Despite these efforts, current methods rely on meticulously crafted graph mining procedures rather than setting graph generation in a learning framework, limiting the inclusion of task graph representations in end-to-end systems.
@@
In this work, we propose an approach to learn task graphs from demonstrations in the form of sequences of key-steps performed by real users in a video while executing a procedure. Given a directed graph represented as an adjacency matrix and a set of key-step sequences, we provide an estimate of the likelihood of observing the set of sequences given the constraints encoded in the graph. We hence formulate task graph learning under the well-understood framework of Maximum Likelihood (ML) estimation, and propose a novel differentiable Task Graph Maximum Likelihood (TGML) loss function which can be naturally plugged into any neural-based architecture for direct optimization of task graph from data. Intuitively, our TGML loss generates positive gradients to strengthen the weights of directed edges $B\to A$ when observing the $<\ldots,A,\ldots,B,\ldots>$ structure, while pushing down the weights of all other edges in a contrastive manner (see Figure 1(b). To evaluate the effectiveness of the proposed framework, we propose two approaches to task graph learning. The first approach, “Direct Optimization (DO)” uses the proposed TGML loss to directly optimize the weights of the adjacency matrix, which constitute the only parameters of the model. The output of the optimization procedure is the learned graph. The second approach, termed Task Graph Transformer (TGT) is a feature-based model which uses a transformer encoder and a relation head to predict the adjacency matrix from either text or video key-step embeddings.
@@
The contributions of this work are: 1) We introduce a novel framework for learning task graphs from action sequences, which relies on maximum likelihood estimation to provide a differentiable loss function which can be included in end-to-end models and optimized with gradient descent; 2) We propose two approaches to task graph learning based on direct optimization of the adjacency matrix and processing key-step text or video embeddings, which offer significant improvements over previous methods in task graph generation and shows emerging video understanding abilities; 3) We showcase the usefulness of task graphs in general, and the learned graph-based representations in particular, on the downstream task of online mistake detection from video, where we improve over competitors. The code to replicate the experiments is available at https://github.com/fpv-iplab/Differentiable-Task-Graph-Learning.
@@ 3.1 Task Graph Maximum Likelihood Learning Framework
Modeling Sequence Likelihood for an Unweighted Graph  Let us consider the special case of an unweighted graph, i.e., $\bar{Z}\in\{0,1\}^{n\times n}$. We wish to estimate $P(y|Z)$, the likelihood of the generic sequence $y\in\mathcal{Y}$ given graph $Z$. Formally, let $Y_{t}$ be the random variable related to the event “key-step $K_{y_{t}}$ appears at position $t$ in sequence $y$”. We can factorize the conditional probability $P(y|Z)$ as:
@@
Modeling Sequence Likelihood for a Weighted Graph  In practice, to enable gradient-based learning, we consider the general case of a continuous adjacency matrix $Z\in[0,1]^{n\times n}$. We generalize the concept of “possible cases” discussed in the previous section with the concept of “feasibility of sampling a given key-step $K_{i}$, having observed a set of key-steps $K_{\mathcal{J}}$, given graph $Z$”, which we define as the sum of all weights of edges between observed key-steps $K_{\mathcal{J}}$ and $K_{i}$: $f(K_{i}|K_{\mathcal{J}},Z)=\sum_{j\in\mathcal{J}}Z_{ij}$. Intuitively, if key-step $k_{i}$ has many satisfied pre-conditions, we are more likely to sample it as the next key-step. We hence define $P(K_{i}|K_{\mathcal{J}},Z)$ as “the ratio of the feasibility of sampling $K_{i}$ to the sum of the feasibilities of sampling any unobserved key-step”:
@@
Figure 2: Given a sequence $<S,A,B,D,C,E>$, and a graph $G$ with adjacency matrix $Z$, our goal is to estimate the likelihood $P(<S,A,B,C,D,E>|Z)$, which can be done by factorizing the expression into simpler terms. The figure shows an example of computation of probability $P(D|S,A,B,Z)$ as the ratio of the “feasibility of sampling key-step D, having observed key-steps S, A, and B” to the sum of all feasibility scores for unobserved symbols. Feasibility values are computed by summing weights of edges $D\to X$ for all observed key-steps $X$.
Figure 2 illustrates the computation of the likelihood in Eq. 3. Plugging Eq. (3) into Eq. (1), we can estimate the likelihood of a sequence $y$ given graph $Z$ as:
@@
where $\beta$ is a hyper-parameter. We refer to Eq. (6) as the Task Graph Maximum Likelihood (TGML) loss function. Since Eq. (6) is differentiable with respect to all $Z_{ij}$ values, we can learn the adjacency matrix $Z$ by minimizing the loss with gradient descent to find the estimated graph $\hat{Z}=\arg_{Z}\max\mathcal{L}(\mathcal{Y},Z)$. Eq. (6) works as a contrastive loss in which the first logarithmic term aims to maximize, at every step $t$ of each input sequence, the weights $Z_{y_{t}j}$ of edges $K_{y_{t}}\to K_{j}$ going from the current key-step $K_{y_{t}}$ to all previously observed key-steps $K_{j}$, while the second logarithmic term (contrastive term) aims to minimize the weights of edges $K_{h}\to K_{j}$ between key-steps yet to appear $K_{h}$ and already observed key-steps $K_{j}$. The hyper-parameter $\beta$ regulates the influence of the summation in the contrastive term which, including many more addends, can dominate gradient updates.
@@ 3.2 Models
Figure 3: Our Task Graph Transformer (TGT) takes as input either D-dimensional text embeddings or video embeddings extracted from key-step segments. Learnable $S$ and $E$ embeddings are also included. Key-step embeddings are processed with a transformer encoder and regularized with a distinctiveness cross-entropy to prevent representation collapse. Output embeddings are processed with our relation head which concatenates vectors across all $(n+2)^{2}$ possible node pairs, thus obtaining $(n+2)\times(n+2)\times 2D$ relation vectors. These are processed by a relation transformer which progressively maps relation vectors to an $(n+2)\times(n+2)$ adjacency matrix. The model is supervised with input sequences using our proposed Task Graph Maximum Likelihood (TGML) loss.
@@ 4.1 Graph Generation
Table 1: Task graph generation results on CaptainCook4D. Best results are in bold, second best results are underlined, best results among competitors are highlighted. Confidence interval bounds computed at $95\%$ conf. for $5$ runs.
@@
TGT-text (Ours) 71.7 ±$\pm 3.6$ 72.9 ±$\pm 2.8$ 72.1 ±$\pm 3.2$
DO (Ours) 86.4 ±$\pm 0.8$ 89.7 ±$\pm 0.5$ 87.8 ±$\pm 0.6$
@@
Random 50.00 50.00
TGT-video 63.11 62.19
Improvement +13.11 +12.19
@@
Graph Generation Results  Results in Table 1 highlight the complexity of the task, with classic approaches based on inductive logic, such as MSGI, achieving poor performance ($12.8$ $F_{1}$), language models and count-based statistics reconstructing only basic elements of the graph ($55.0$ and $61.0$ $F_{1}$ for LLM and Count-Based respectively), and even more recent methods based on inductive logic and heuristics only partially predicting the graph ($71.1$ $F_{1}$ of $MSG^{2}$). The proposed Direct Optimization (DO) approach outperforms all other methods, achieving the highest scores across all measures, with improvements in the $[+15.5,+18.1]$ range with respect to the best competitor $MSG^{2}$. This result highlights the effectiveness of the proposed framework to learn task graph representations from key-step sequences, especially considering the simplicity of the DO method, which performs gradient descent directly on the adjacency matrix. We obtain a slightly higher recall as compared to the precision ($89.7$ vs $86.4$), showing that our approach tends to retrieve most ground truth edges, while hallucinating some pre-conditions, probably due to the dataset being unbalanced towards the most common ways of completing a procedure. Second best results are consistently obtained by our feature-based TGT approach, showing the generality of our learning framework and the potential of integrating it into complex neural architectures. Tight confidence intervals for DO highlight the stability of the proposed loss. The lower performance of TGT, as compared to DO, may be due to the relatively small size of the dataset, which makes it hard for complex architecture to generalize.
Video Understanding Results  Table 2 reports the performance of TGT trained on videos on two fundamental video understanding tasks [42] of pairwise clip ordering and future prediction. [^5] See the supplementary material for more details. For pairwise ordering, we feed our TGT model with video embeddings of two clips and sort them according to the predicted adjacency matrix, placing first the clip identified as a pre-condition. For future predictions, given an anchor clip, we have to choose which among two other clips is the correct future. In this case, we feed the model with the three clips and select as future the clip with the smallest $anchor\to clip$ edge weight (future clips are not likely pre-conditions). Despite TGT not being explicitly trained for pairwise ordering and future predictions, it exhibits promising emerging video understanding abilities, surpassing the random baseline.
@@ 4.2 Online Mistake Detection
Results  Results in Table 3 highlight the usefulness of the learned task graphs for downstream applications. The proposed DO method achieves significant gains over prior art with improvements of $+19.8$ and $+7.5$ in average $F_{1}$ score on Assembly101 and EPIC-Tent respectively when ground truth action sequences are considered to make predictions. While TGT is the second-best performer on Assembly101, it obtains best results on EPIC-Tent ($64.1$ vs $58.3$ in average $F_{1}$ score). This is due to the nature of action annotations in the two datasets. Indeed, while key-step names are informative in EPIC-Tent (e.g., “Place Vent Cover”, “Open Stake Bag”, or “Spread Tent”), they are less distinctive in Assembly101 (e.g., “attach cabin”, “attach interior”, or “screw chassis”). This highlights the flexibility of the proposed learning framework which can work in purely abstract, symbolic settings, with the DO approach, but can also leverage semantics with TGT when beneficial. Interestingly, the second best performers are graph-based approaches, with $MSG^{2}$ achieving an average $F_{1}$ of $56.1$ on Assembly101 and the simple Count-Based approach obtaining an average $F_{1}$ score of $56.6$ on EPIC-Tent. In contrast, PREGO obtains average $F_{1}$ scores of $39.4$ and $32.1$ on Assembly101 and EPIC-Tent respectively, suggesting the potential of explicit graph-based representations for mistake detection, versus the implicit one of PREGO. Breaking down performance into correct and mistake $F_{1}$ scores reveal some degree of unbalance of our approaches and the main competitor $MSG^{2}$ towards identifying correct actions rather than mistakes. This suggests that the related graph-based representations tend to detect some spurious pre-conditions, probably due to the limited demonstrations included in the videos, while the implicit PREGO model exhibits a skew with respect to mistakes. Further breaking down $F_{1}$ scores into related precision and recall values highlights that the main failure modes are due to large imbalances between precision and recall. For instance, the Count-Based method achieves a precision of only $5.1$ with a recall of $86.2$ in predicting correct segments on Assembly101. In contrast, the proposed approach obtains balanced precision and recall values in detecting correct segments in Assembly101 ($98.2$/$83.4$) and EPIC-Tent ($94.8$/$92.4$), and detecting mistakes in EPIC-Tent ($33.2$/$35.7$), while the prediction of mistakes on Assembly101 is more skewed ($46.7$/$90.4$). Results based on action sequences predicted from video (bottom part of Table 3) highlight the challenging nature of the task, when considering noisy action sequences. While the explicit task graph representation may not accurately reflect the predicted noisy action sequences, we still observe improvements over previous approaches of $+7.3$ and $+1.3$ in average $F_{1}$ score in Assembly101 and EPIC-Tent. Remarkably, best competitors are still graph-based methods, such as $MSG^{2}$ and the Count-Based approach, with significant improvements over the implicit representation of the PREGO model ($32.5$ average $F_{1}$ versus $53.5$ of the proposed DO model). Also in this case, we observe that graph-based methods tend to be skewed towards detecting correct action sequences. In this regard, our TGT model only achieves $38.2$ mistake $F_{1}$ score, a drop in $5.5$ points over the best performer, the Count-Based method, which, on the other hand, only achieves an $F_{1}$ score of $2.5$ when predicting correct segments.
@@
Assembly101 EPIC-Tent

### v3

@@ Abstract
Procedural activities are sequences of key-steps aimed at achieving specific goals. They are crucial to build intelligent agents able to assist users effectively. In this context, task graphs have emerged as a human-understandable representation of procedural activities, encoding a partial ordering over the key-steps. While previous works generally relied on hand-crafted procedures to extract task graphs from videos, in this paper, we propose an approach based on direct maximum likelihood optimization of edges’ weights, which allows gradient-based learning of task graphs and can be naturally plugged into neural network architectures. Experiments on the CaptainCook4D dataset demonstrate the ability of our approach to predict accurate task graphs from the observation of action sequences, with an improvement of +16.7% over previous approaches. Owing to the differentiability of the proposed framework, we also introduce a feature-based approach, aiming to predict task graphs from key-step textual or video embeddings, for which we observe emerging video understanding abilities. Task graphs learned with our approach are also shown to significantly enhance online mistake detection in procedural egocentric videos, achieving notable gains of +19.8% and +7.5% on the Assembly101-O and EPIC-Tent-O datasets. Code for replicating the experiments is available at https://github.com/fpv-iplab/Differentiable-Task-Graph-Learning.
@@ 1 Introduction
Previous approaches focused on directly tackling tasks requiring procedural knowledge such as action anticipation [16, 14, 34] and mistake detection [37, 13, 7, 41, 15] without developing explicit representations of the procedure. Other works proposed neural models able to develop implicit representations of the procedure by learning how to recover missing actions [43, 28], discover key-steps [11, 4, 5], or grounding them to video [10, 25]. A different approach [3, 10, 18] consists in representing the structure of a procedure in the form of a task graph, i.e., a Directed Acyclic Graph (DAG) in which nodes represent key-steps, and directed edges impose a partial ordering over key-steps, encoding dependencies between them (see Figure 1(a)). [^1] See the supplementary material for more details. Graphs provide an explicit representation which is readily interpretable by humans and easy to incorporate in downstream tasks such as detecting mistakes or validating the execution of a procedure. While graphs have been historically used to represent constraints in complex tasks and design optimal sub-tasks scheduling [38], graph-based representations mined from videos [3], key-step sequences [39, 20] or external knowledge bases [44] have only recently emerged as a powerful representation of procedural activities able to support downstream tasks such as key-step recognition or forecasting [3, 44]. Despite these efforts, current methods rely on meticulously crafted graph mining procedures rather than setting graph generation in a learning framework, limiting the inclusion of task graph representations in end-to-end systems.
@@
In this work, we propose a novel approach to learn task graphs from demonstrations in the form of sequences of key-steps performed by real users in a video while executing a procedure. Given a directed graph represented as an adjacency matrix and a set of key-step sequences, we provide an estimate of the likelihood of observing the set of sequences given the constraints encoded in the graph. We hence formulate task graph learning under the well-understood framework of Maximum Likelihood (ML) estimation, and propose a novel differentiable Task Graph Maximum Likelihood (TGML) loss function which can be naturally plugged into any neural-based architecture for direct optimization of task graph from data. Intuitively, our TGML loss generates positive gradients to strengthen the weights of directed edges $B\to A$ when observing the $<\ldots,A,\ldots,B,\ldots>$ structure, while pushing down the weights of all other edges in a contrastive manner (see Figure 1(b)). To evaluate the effectiveness of the proposed framework, we propose two approaches to task graph learning. The first approach, called “Direct Optimization (DO)”, uses the proposed TGML loss to directly optimize the weights of the adjacency matrix, which constitute the only parameters of the model. The output of the optimization procedure is the learned graph. The second approach, termed Task Graph Transformer (TGT) is a feature-based model which uses a transformer encoder and a relation head to predict the adjacency matrix from either text or video key-step embeddings.
@@
The contributions of this work are the following: 1) We introduce a novel framework for learning task graphs from action sequences, which relies on maximum likelihood estimation to provide a differentiable loss function which can be included in end-to-end models and optimized with gradient descent; 2) We propose two approaches to task graph learning based on direct optimization of the adjacency matrix and processing key-step text or video embeddings, which offer significant improvements over previous methods in task graph generation and shows emerging video understanding abilities; 3) We showcase the usefulness of task graphs in general, and the learned graph-based representations in particular, on the downstream task of online mistake detection from video, where we improve over competitors. The code to replicate the experiments is available at https://github.com/fpv-iplab/Differentiable-Task-Graph-Learning.
@@ 3.1 Task Graph Maximum Likelihood Learning Framework
Modeling Sequence Likelihood for an Unweighted Graph  Let us consider the special case of an unweighted graph, i.e., $\bar{Z}\in\{0,1\}^{(n+2)\times(n+2)}$. We wish to estimate $P(y|Z)$, the likelihood of the generic sequence $y\in\mathcal{Y}$ given graph $Z$. Formally, let $Y_{t}$ be the random variable related to the event “key-step $K_{y_{t}}$ appears at position $t$ in sequence $y$”. We can factorize the conditional probability $P(y|Z)$ as:
@@
Modeling Sequence Likelihood for a Weighted Graph  To enable gradient-based learning, we consider the general case of a continuous adjacency matrix $Z\in[0,1]^{(n+2)\times(n+2)}$. We generalize the concept of “possible cases” discussed in the previous section with the concept of “feasibility of sampling a given key-step $K_{i}$, having observed a set of key-steps $K_{\mathcal{J}}$, given graph $Z$”, which we define as the sum of all weights of edges between observed key-steps $K_{\mathcal{J}}$ and $K_{i}$: $f(K_{i}|K_{\mathcal{J}},Z)=\sum_{j\in\mathcal{J}}Z_{ij}$. Intuitively, if key-step $k_{i}$ has many satisfied pre-conditions, we are more likely to sample it as the next key-step. We hence define $P(K_{i}|K_{\mathcal{J}},Z)$ as “the ratio of the feasibility of sampling $K_{i}$ to the sum of the feasibilities of sampling any unobserved key-step”:
@@
Figure 2: Given a sequence $<S,A,B,D,C,E>$, and a graph $G$ with adjacency matrix $Z$, our goal is to estimate the likelihood $P(<S,A,B,D,C,E>|Z)$, which can be done by factorizing the expression into simpler terms. The figure shows an example of computation of probability $P(D|S,A,B,Z)$ as the ratio of the “feasibility of sampling key-step D, having observed key-steps S, A, and B” to the sum of all feasibility scores for unobserved symbols. Feasibility values are computed by summing weights of edges $D\to X$ for all observed key-steps $X$.
Figure 2 illustrates the computation of the likelihood in Eq. (3). Plugging Eq. (3) into Eq. (1), we can estimate the likelihood of a sequence $y$ given graph $Z$ as:
@@
where $\beta$ is a hyper-parameter. We refer to Eq. (6) as the Task Graph Maximum Likelihood (TGML) loss function. Since Eq. (6) is differentiable with respect to all $Z_{ij}$ values, we can learn the adjacency matrix $Z$ by minimizing the loss with gradient descent to find the estimated graph $\hat{Z}=\arg_{Z}\max\mathcal{L}(\mathcal{Y},Z)$. Eq. (6) works as a contrastive loss in which the first logarithmic term aims to maximize, at every step $t$ of each input sequence, the weights $Z_{y_{t}j}$ of edges $K_{y_{t}}\to K_{j}$ going from the current key-step $K_{y_{t}}$ to all previously observed key-steps $K_{j}$, while the second logarithmic term (contrastive term) aims to minimize the weights of edges $K_{h}\to K_{j}$ between key-steps yet to appear $K_{h}$ and already observed key-steps $K_{j}$. The hyper-parameter $\beta$ regulates the influence of the summation in the contrastive term which, including many more addends, can dominate gradient updates. As in other contrastive learning frameworks [29, 33], our approach only includes positives and negatives and it does not explicitly consider anchor examples.
@@ 3.2 Models
Figure 3: Our Task Graph Transformer (TGT) takes as input either $D$-dimensional text embeddings extracted from key-step names or video embeddings extracted from key-step segments. In both cases, we extract features with a pre-trained EgoVLPv2 model. For video embeddings, multiple embeddings can refer to the same action, so we randomly select one for each key-step (RS blocks). Learnable start (S) and end (E) embeddings are also included. Key-step embeddings are processed using a transformer encoder and regularized with a distinctiveness cross-entropy to prevent representation collapse. The output embeddings are processed by our relation head, which concatenates vectors across all $(n+2)^{2}$ possible node pairs, producing $(n+2)\times(n+2)\times 2D$ relation vectors. These vectors are then processed by a relation transformer, which progressively maps them to an $(n+2)\times(n+2)$ adjacency matrix. The model is supervised with input sequences using our proposed Task Graph Maximum Likelihood (TGML) loss.
@@ 4.1 Graph Generation
Table 1: Task graph generation results on CaptainCook4D. Best results are in bold, second best results are underlined, best results among competitors are highlighted. Confidence interval bounds computed at $90\%$ conf. for $5$ runs.
@@
TGT-text (Ours) 79.9 ±$\pm 8.8$ 81.9 ±$\pm 6.9$ 80.8 ±$\pm 8.0$
DO (Ours) 86.4 ±$\pm 1.5$ 89.7 ±$\pm 1.5$ 87.8 ±$\pm 1.5$
@@
Random 50.0 50.0
TGT-video 77.3 74.3
Improvement +27.3 +24.3
@@
Graph Generation Results  Results in Table 1 highlight the complexity of the task, with classic approaches based on inductive logic, such as MSGI, achieving poor performance ($12.8$ $F_{1}$), language models and count-based statistics reconstructing only basic elements of the graph ($55.0$ and $60.6$ $F_{1}$ for LLM and Count-Based respectively), and even more recent methods based on inductive logic and heuristics only partially predicting the graph ($71.1$ $F_{1}$ of $MSG^{2}$). The proposed Direct Optimization (DO) approach outperforms all other methods, achieving the highest scores across all measures, with improvements in the $[+15.5,+18.1]$ range with respect to the best competitor $MSG^{2}$. This result highlights the effectiveness of the proposed framework to learn task graph representations from key-step sequences, especially considering the simplicity of the DO method, which performs gradient descent directly on the adjacency matrix. We obtain a slightly higher recall as compared to the precision ($89.7$ vs $86.4$), showing that our approach tends to retrieve most ground truth edges, while hallucinating some pre-conditions, probably due to the dataset being unbalanced towards the most common ways of completing a procedure. Second best results are consistently obtained by our feature-based TGT approach, showing the generality of our learning framework and the potential of integrating it into complex neural architectures. Tight confidence intervals for DO highlight the stability of the proposed loss. The lower performance of TGT, as compared to DO, may be due to the relatively small size of the dataset, which makes it hard for complex architecture to generalize.
Video Understanding Results  Table 2 reports the performance of TGT trained on videos on two fundamental video understanding tasks [46] of pairwise clip ordering and future prediction. [^6] See the supplementary material for more details. For pairwise ordering, we feed our TGT model with video embeddings of two clips and sort them according to the predicted adjacency matrix, placing first the clip identified as a pre-condition. For future predictions, given an anchor clip, we have to choose which among two other clips is the correct future. Despite TGT not being explicitly trained for pairwise ordering and future predictions, it exhibits emerging video understanding abilities, surpassing the random baseline.
@@ 4.2 Online Mistake Detection
Results  Results in Table 3 highlight the usefulness of the learned task graphs for downstream applications. The proposed DO method achieves significant gains over prior art with improvements of $+19.8$ and $+7.5$ in average $F_{1}$ score on Assembly101-O and EPIC-Tent-O respectively when ground truth action sequences are considered to make predictions. While TGT is the second-best performer on Assembly101-O, it obtains best results on EPIC-Tent-O ($64.1$ vs $58.3$ in average $F_{1}$ score). This is due to the nature of action annotations in the two datasets. Indeed, while key-step names are informative in EPIC-Tent (e.g., “Place Vent Cover”, “Open Stake Bag”, or “Spread Tent”), they are less distinctive in Assembly101 (e.g., “attach cabin”, “attach interior”, or “screw chassis”). This highlights the flexibility of the proposed learning framework which can work in purely abstract, symbolic settings, with the DO approach, but can also leverage semantics with TGT when beneficial. Interestingly, the second best performers are graph-based approaches, with $MSG^{2}$ achieving an average $F_{1}$ of $56.1$ on Assembly101-O and the simple Count-Based approach obtaining an average $F_{1}$ score of $56.6$ on EPIC-Tent-O. In contrast, PREGO obtains average $F_{1}$ scores of $39.4$ and $32.1$ on Assembly101-O and EPIC-Tent-O respectively, suggesting the potential of explicit graph-based representations for mistake detection, versus the implicit one of PREGO. Breaking down performance into correct and mistake $F_{1}$ scores reveal some degree of unbalance of our approaches and the main competitor $MSG^{2}$ towards identifying correct actions rather than mistakes. This suggests that the related graph-based representations tend to detect some spurious pre-conditions, probably due to the limited demonstrations included in the videos, while the implicit PREGO model exhibits a skew with respect to mistakes. Further breaking down $F_{1}$ scores into related precision and recall values highlights that the main failure modes are due to large imbalances between precision and recall. For instance, the Count-Based method achieves a precision of only $4.8$ with a recall of $85.7$ in predicting correct segments on Assembly101-O. In contrast, the proposed approach obtains balanced precision and recall values in detecting correct segments in Assembly101-O ($98.2$/$83.4$) and EPIC-Tent-O ($94.1$/$93.5$), and detecting mistakes in EPIC-Tent-O ($33.3$/$35.7$), while the prediction of mistakes on Assembly101-O is more skewed ($46.7$/$90.4$). Results based on action sequences predicted from video (bottom part of Table 3) highlight the challenging nature of the task when considering noisy action sequences (see Figure 4). While the explicit task graph representation may not accurately reflect the predicted noisy action sequences, we still observe improvements over previous approaches of $+7.3$ and $+1.3$ in average $F_{1}$ score in Assembly101-O and EPIC-Tent-O. Remarkably, best competitors are still graph-based methods, such as $MSG^{2}$ and the Count-Based approach, with significant improvements over the implicit representation of the PREGO model ($32.5$ average $F_{1}$ versus $53.5$ of the proposed DO model). Also, in this case, we observe that graph-based methods tend to be skewed towards detecting correct action sequences. In this regard, our TGT model only achieves $38.2$ in mistake $F_{1}$ score, a drop in $5.7$ points over the best performer, the Count-Based method, which, on the other hand, only achieves an $F_{1}$ score of $2.6$ when predicting correct segments.
@@
Assembly101-O EPIC-Tent-O